OpenStudy (anonymous):

Find the value of the variable that would make the equation true. 2/3 x =8

jigglypuff314 (jigglypuff314):

multiply both sides by 3 then divide both sides by 2

OpenStudy (shamil98):

\[\large \frac{ 2 }{ 3 } x = 8\] multiply both sides by the reciprocal. \[\large \frac{ \cancel 3 }{ \cancel 2 } \times \frac{ \cancel 2 }{ \cancel 3 } x = 8 \times \frac{ 3 }{ 2 }\] \[\large x = \frac{ 24 }{ 2 } = 12\]